OpenStudy (amber15):

pair of markings , a set distance apart are made on highways so that police can detect drivers exceeding over the speed limit . over a fixed distance , the speed R varies inversely with time T . For one particular pair of markings , R is 50 mph when T is 7 seconds . find the speed of a car that travels the given distance in 5 seconds. please help

jimthompson5910 (jim_thompson5910):

R varies inversely with time T means R = k/T for some constant k

jimthompson5910 (jim_thompson5910):

R is 50 mph when T is 7 seconds so plug in R = 50 and T = 7 and solve for k R = k/T 50 = k/7 k = ???

OpenStudy (amber15):

350

jimthompson5910 (jim_thompson5910):

so the equation is R = 350/T

jimthompson5910 (jim_thompson5910):

"find the speed of a car that travels the given distance in 5 seconds." now plug in T = 5 to find R

OpenStudy (amber15):

okay

jimthompson5910 (jim_thompson5910):

what do you get?

OpenStudy (amber15):

355

OpenStudy (amber15):

oops mistake

jimthompson5910 (jim_thompson5910):

R = 350/T R = 350/5 ... replace T with 5 R = ??

OpenStudy (amber15):

70

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (amber15):

Thank you do to i'm not good on these problems

jimthompson5910 (jim_thompson5910):

you're welcome